OpenStudy (anonymous):

If $7300 is invested in a savings account for which interest is compounded monthly, and if the $7300 turns into $8600 in 2 years, what is the interest rate of the savings account? A.16.50% B.8.22% C.4.10% D.5.48%

OpenStudy (kropot72):

The equation needed is \[8600=7300(1+\frac{r}{12})^{24}\] where r is the annual interest rate as a decimal.

OpenStudy (primeralph):

24 or 2?

OpenStudy (kropot72):

The expression in brackets is raised to the power of 24, the reason being that the interest is compounded monthly and there are 24 months in two years. Note that the annual interest rate is divided by 12 to give a monthly rate.

OpenStudy (primeralph):

yeah, didn't see that.

OpenStudy (kropot72):

You're welcome :)

OpenStudy (whpalmer4):

The formula is sometimes written\[FV = PV(1+\frac{r}{n})^{nt}\]where \(n\) is the number of compounding periods per year, and \(t\) the number of years. You can easily see that if \(n=1\) it reduces to the annual compounding case.